A Complete Bibliography of Publications in Applied Mathematics and Computation: 2025–2029

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

09 November 2024 Version 1.03

Title word cross-reference

1 [JSZ25]. 1n [AA25]. 3 [MR25]. 4 [LZ25]. α [GCV25]. B [CCLL25]. M [AA25].

-equivalent [CCLL25]. -factorization [JSZ25]. -fractal [GCV25]. -independence [LZ25]. -Wright [AA25].

1 [Ano25a, Ano25b]. 15 [Ano25c, Ano25d].

2025 [Ano25a, Ano25c, Ano25b, Ano25d].

395 [AS25].

accounting [LL25]. Acoustics [ALC^{+}25]. action [MMNP25]. active [CGH^{+}25]. actuator [BAT25, HWC25, SNP25, ZZ25a].

Adaptive [ZZ25b, BAT25, LHZL25, LWS^{+}25, ZLA25, ZLCL25]. advection [PP25, dOKA25]. advection-diffusion [PP25]. affine [IJM25]. against [SNP25]. agent [LHZL25, ZZ25a]. algorithm [MCHB25, Pal25]. algorithms [LTH^{+}25]. Allee [CMG25]. Allen [OGPFVV25]. allocation [RS25]. almost [CMG25]. Analysis [FBK25, GCV25, ALC^{+}25, BSG^{+}25, CWWT25, RPS^{+}25]. Analytically [Ruj25]. aperiodic [WSX^{+}25]. Appl [AS25]. application [ZJX25]. applications [Liu25]. applied [TNL25]. approach [DAG25]. approximation [FLLR25, IJM25]. arbitrary [CWWT25]. artificial [AS21, AS25]. assets [XZN^{+}25]. associative [SNP25]. asymmetric [GLHS25]. asynchronous [MCH25]. Attack [MCH25]. Attack-compensated [MCH25]. attacks [BAT25, ZLCL25]. attitude [ZLCL25]. avoidance [ZLA25].

B [PP25]. B-splines [PP25]. background [RS25]. Barrier [ZLA25]. Barrier-function [ZLA25]. based [CWWT25, DWL^{+}25, HWC25, LHZL25, LWS^{+}25, LM25, PP25, Pal25, QX25, WSX^{+}25, ZWQC25, ZLA25]. basis [TNL25]. bead [MBF25]. bead-rod [MBF25]. bidirectional [SNP25]. bilinear [IJM25]. biological [LM25]. bipartite [SHY25, YT25]. blind [IAK25]. blood [RPS^{+}25]. Board [Ano25e, Ano25f, Ano25g, Ano25h]. boundary [ALC^{+}25, GCV25, OGPFVV25]. branch [BVZ25]. Brownian [MBF25]. burning [LSQ25].

Cahn [OGPFVV25]. capped [Ruj25]. capped/floored [Ruj25]. case [Ruj25]. cell [LM25]. center [LM25]. center-based [LM25]. chains [MBF25, ZW25]. Chebyshev [GKM25]. class [BAT25]. classes [ZW25]. codes [Kov25]. coefficients [CWWT25]. Cohen [ZN25]. collision [ZLA25]. collocation [PP25]. combination [Liu25]. communication [WSX^{+}25, ZZ25a]. compact [MR25]. compensated [MCH25]. competitive [FBK25]. complete [JSZ25, LYM25]. Component [ZLW25]. components [BQS25]. Compt [AS25]. computation [BQS25]. conditions [ALC^{+}25, GCV25, KL25, Liu25, OGPFVV25]. configuration [MBF25]. conjugate [CFS25]. connectivity [ZLW25]. consensus [GPL25, ZZ25a]. conservation [KBM25, TNL25]. considering [WSX^{+}25]. constrained [GPL25]. constraint [HWC25]. constraints [LWS^{+}25]. Contagion [KK25a]. control [BAT25, CGH^{+}25, GPL25, HWC25, IJM25, KNKR25, LHZL25, LWS^{+}25, MCH25, QX25, SNP25, ZWQC25, ZLA25, ZZ25a, ZLCL25]. control-affine [IJM25]. convergence [BSG^{+}25]. convexities [ADPS25]. cooperation [GLHS25, LL25, YYW^{+}25]. Corrigendum [AS25]. cospectral [LSWZ25]. cover [LCL25, ZHL25]. covers [QM25]. cross [FBK25]. cross-immunity [FBK25]. cycle [LCL25]. cyclic [ZW25].

D [MR25]. Data [IJM25, WSX^{+}25]. Data-driven [IJM25]. DDSIIM [CWWT25]. dead [LHZL25]. dead-zone [LHZL25]. deblurring [IAK25]. deception [ZLCL25]. decision [ZJX25]. decision-making [ZJX25]. degenerate [KK25b]. delay [WSX^{+}25]. delayed [CCLL25, SZY25, SNP25, WSX^{+}25]. delays [ZN25]. dependent [BQS25, CCLL25]. derivative [LT25]. derivatives [AS21, AS25]. described [XZN^{+}25]. description [MMNP25]. determinant [ZJX25]. deterministic [DAG25]. diameter [LSWZ25]. difference [OGPFVV25]. differential [GPL25, XZN^{+}25]. diffusion [PP25, Ruj25, dOKA25]. dilemma [YYW^{+}25]. dilute [MBF25]. dimensional [LJ25]. directions [QX25]. Dirichlet [ALC^{+}25]. discontinuous [CWWT25]. discrete [IJM25, Ruj25]. discrete-time [IJM25]. disease [FBK25]. Disjoint [QM25, LCL25]. Distance [ZW25]. distributed [ZZ25a, ZZ25b]. distribution [MP25]. Disturbance [LWS^{+}25]. disturbances [CGH^{+}25]. domains [CWWT25, TKM25]. domatic [GVW25]. domination [GVW25]. driven [GPL25, IJM25, Ruj25]. Dynamic [GPL25, ZLCL25, HWC25]. dynamical [CGH^{+}25]. dynamics [CMG25, Che25, IJM25, LDT^{+}25].

Editorial

[Ano25e, Ano25f, Ano25g, Ano25h]. effect [CMG25]. effects [FBK25]. Efficient [LT25, TKM25]. eigenvalue [LYM25, MP25]. element [LT25, LJ25, Tan25, TKM25]. elliptic [CWWT25]. energies [ZW25]. enhanced [LT25]. Enhancement [LDT^{+}25]. entropy [CLLC25]. Epidemic [FBK25]. epidemics [MMNP25]. equality [LL25]. equation [GKM25, LTH^{+}25, PP25]. equations [ALC^{+}25, FLLR25, MR25, OGPFVV25, Tan25, TKM25, XZN^{+}25]. equienergetic [LSWZ25]. equivalent [CCLL25, Liu25]. Erdos [MMNP25]. error [CWWT25, CLLC25]. errors [ZJX25]. estimates [TNL25]. Estimation [XZN^{+}25, ZWQC25, SZY25, dOKA25]. Estimation-based [ZWQC25]. Euler [TKM25]. Event [QX25, ZZ25a, GPL25, HWC25, LHZL25, ZWQC25, ZLCL25]. event-driven [GPL25]. Event-triggered [QX25, ZZ25a, HWC25, LHZL25, ZWQC25, ZLCL25]. Evolution [GLHS25, DWL^{+}25]. evolutionary [LL25]. expansion [BSG^{+}25]. explicit [dOKA25]. Exploring [MMNP25, YT25]. Exponential [CCLL25, MR25, ZN25]. extend [CLLC25]. extinction [CMG25]. extrema [LZ25].

factorization [JSZ25]. failure [SNP25]. fairness [DWL^{+}25]. fast [Pal25, Tan25]. Fault [KNKR25, CGH^{+}25]. Fault-tolerant [KNKR25, CGH^{+}25]. faults [BAT25, CGH^{+}25, HWC25, ZZ25a]. February [Ano25a, Ano25c]. feedback [BAT25, MCH25, SNP25]. FFT [CWWT25]. FFT-based [CWWT25]. fiducial [CLLC25]. field [MMNP25]. fields [MBF25]. filter [CLLC25]. Finite [IJM25, CF25, HWC25, LT25, LJ25, OGPFVV25, Tan25, ZJX25]. finite-difference [OGPFVV25]. finite-time [HWC25]. first [GKM25]. Fixed [CGH^{+}25, QX25, ZWQC25]. Fixed-time [CGH^{+}25, QX25, ZWQC25]. flight [KNKR25]. floored [Ruj25]. flow [MBF25]. fluid [KBM25]. fluid-poroelastic [KBM25]. forces [LM25]. form [KL25, MR25]. formation [KNKR25]. four [LSWZ25]. fourth [PP25]. fractal [GCV25]. fractional [SHY25, Tan25, XZN^{+}25, ZN25]. fragile [SNP25]. framework [MAMFBM^{+}25]. free [LT25, QX25]. frictions [RS25]. full [LWS^{+}25]. function [AA25, TNL25, ZLA25]. functions [GCV25, MCHB25]. fuzzy [KL25, LWS^{+}25, ZWQC25].

game [YYW^{+}25]. games

[GPL25, LL25, ZJX25]. gasket [GCV25]. gauge [SHY25]. general [ADPS25]. generalized [LZ25, LJ25, ZHL25]. given [LZ25]. global [FLLR25]. glucose [AS21, AS25]. glucose-insulin [AS21]. gradient [CFS25]. graph [ADPS25]. graphical [GPL25]. Graphs [LSQ25, GVW25, LYM25, LZ25, MP25, QM25, YT25]. grid [Tan25]. Grossberg [ZN25]. group [ZJX25]. growth [BDL25]. guidance [YYW^{+}25].

hand [CF25]. harmonic [KK25b]. heterogeneous [ZZ25a]. high [CLLC25, LT25, MR25, ZWQC25, ZLA25]. high-order [CLLC25, LT25, MR25, ZWQC25, ZLA25]. higher [KK25b, LDT^{+}25]. higher-order [KK25b, LDT^{+}25]. Hilbert [Liu25]. Hilbert-type [Liu25]. homogeneous [LJ25]. human [YYW^{+}25]. hyperbolic [CMG25, PP25]. hypergraph [JSZ25]. hypergraphs [MAMFBM^{+}25]. hypersingular [GKM25].

identities [KK25b]. image [IAK25]. immobile [Tan25]. immunity [FBK25]. Impenetrability [ALC^{+}25]. implementation [ZJX25]. implicit [MR25]. Improved [MCHB25]. impulses [CCLL25]. impulsive [ZN25]. independence [LZ25]. indices [ZHL25]. inequality [KL25, WSX^{+}25]. infinite [CF25]. infinitesimal [CF25]. influence [DAG25]. injection [BAT25]. input [QX25, ZZ25b]. insulin [AS21, AS25]. insurable [RS25]. Integral [dOKA25, FLLR25, GKM25, Liu25, WSX^{+}25]. integration [LM25]. interaction [KBM25]. interactions [LDT^{+}25]. interface [CWWT25, LJ25]. intermittent [CGH^{+}25]. investment [RS25]. iteration [ADPS25].

January [Ano25b, Ano25d]. Jarratt [BSG^{+}25]. Jarratt-type [BSG^{+}25]. jump [Ruj25]. jump-diffusion [Ruj25].

Kalman [CLLC25]. kernel [Liu25]. kind [FLLR25, GKM25]. kinetics [AS21, AS25]. Koopman [IJM25].

labelling [BVZ25]. Laplacian [LYM25, MP25, ZW25]. law [LTH^{+}25, MMNP25]. laws [TNL25]. learning [LHZL25]. learning-based [LHZL25]. least [LYM25]. Level [IAK25, MR25]. Lexicographic [AS21, AS25]. limits [MMNP25]. Linear [CF25, CMG25, KK25a, KL25, MR25, Tan25, TNL25, ZZ25b, dOKA25]. linear-hyperbolic [CMG25]. Local [LTH^{+}25, Pal25]. Lotka [Che25]. Lur’e [HWC25].

maintains [YYW^{+}25]. making [ZJX25]. manifolds [CFS25]. Mass [KBM25, MMNP25]. matching [ZHL25]. Math [AS25]. matrix [KL25, WSX^{+}25]. matrix-separation-based [WSX^{+}25]. mean [MMNP25]. mechanical [LM25]. mechanism [Che25, DWL^{+}25]. memetic [Pal25]. memory [HWC25, SNP25, ZN25]. memory-based [HWC25]. memristive [SZY25, SNP25]. Mental [LL25]. Merton [Ruj25]. metastatic [BDL25]. method [CFS25, CCLL25, FLLR25, IAK25, LT25, PP25, TKM25]. methods [BSG^{+}25, GKM25, Tan25, TNL25]. mimetic [OGPFVV25]. Minimizing [LYM25]. Minimum [CLLC25]. mobile [Tan25]. mobile/immobile [Tan25]. mode [KNKR25]. model [CMG25, Che25, KK25a, KNKR25, RPS^{+}25, Ruj25]. Modelling [DAG25]. models [BDL25, LM25]. most [LSWZ25]. multi [KNKR25, LHZL25, ZZ25a]. multi-agent [LHZL25, ZZ25a]. multi-UAV [KNKR25]. multiagent [GPL25]. Multiple [ZN25, XZN^{+}25]. Multiscale [MBF25]. multispace [Kov25]. multispaces [Kov25]. multistate [BQS25]. multivariate [MCHB25].

nested [KL25]. networks [LCL25, MMNP25, SZY25, SHY25, SNP25, WSX^{+}25, ZLW25, ZN25]. neural [SZY25, SNP25, WSX^{+}25, ZN25]. neutral [YYW^{+}25]. Newton [LT25]. Non [LSWZ25, SNP25, JSZ25, LJ25, Tan25, dOKA25]. Non-cospectral [LSWZ25]. Non-fragile [SNP25]. non-homogeneous [LJ25]. non-linear [Tan25, dOKA25]. non-uniform [JSZ25, Tan25]. nonhomogeneous [Ruj25]. Nonlinear [CFS25, ALC^{+}25, BAT25, FLLR25, IJM25, LWS^{+}25, LTH^{+}25, QX25, RS25, Ruj25, ZWQC25, ZLA25]. nonlinearity [LTH^{+}25]. nonperiodic [OGPFVV25]. nonsmooth [AS21, AS25]. norm [Liu25]. normal [AS21, AS25]. note [JSZ25]. number [ADPS25, LSQ25, LZ25]. numbers [GVW25, KK25b]. Numerical [ALC^{+}25, BDL25, LM25].

observation [Ruj25]. observer [LWS^{+}25, ZZ25b]. observer-based [LWS^{+}25]. operator [IJM25, Liu25]. operators [OGPFVV25]. Optimal [RS25, GPL25, MCHB25]. optimization [CFS25]. options [Ruj25, XZN^{+}25]. order [AA25, CLLC25, KK25b, LT25, LZ25, LDT^{+}25, MR25, PP25, SZY25, Tan25, ZWQC25, ZLA25]. ordering [Pal25]. output [BAT25, MCH25, SNP25]. output-feedback [BAT25, SNP25].

pancreatic [AS21, AS25]. pancyclicity [LCL25]. paper [LDT^{+}25]. parabolic [MR25]. parallel [Pal25]. parameterized [KL25]. parameters [XZN^{+}25]. passivity [SHY25]. path [QM25]. Pattern [Che25]. payoff [LL25]. payoffs [Ruj25]. Perfect [ZHL25]. performance [HWC25]. periodic [OGPFVV25]. persistence [LDT^{+}25]. perspective [LL25]. phenomenon [AA25]. Pinning [SHY25]. point [GCV25]. points [CLLC25]. Poisson [Ruj25]. polymeric [MBF25]. polynomial [PP25]. polyomino [ZW25]. population [CMG25]. populations [LM25, YYW^{+}25]. poroelastic [KBM25]. portfolio [RS25]. position [ADPS25]. poverty [DAG25]. power [LTH^{+}25]. Practical [HWC25]. predefined [ZLCL25]. predefined-time [ZLCL25]. predictive [KNKR25]. predictor [IJM25]. presence [BAT25]. preserving [LTH^{+}25]. prevalence [DAG25]. pricing [Ruj25]. prisoner [YYW^{+}25]. probability [KK25a, MCH25]. problem [Pal25]. problems [CWWT25, IAK25, LJ25, dOKA25]. process [Ruj25]. programming [CF25]. projection [GKM25]. promotes [DWL^{+}25]. promotion [LL25]. prostitution [DAG25].

quantization [MCHB25]. quantized [HWC25]. quarantine [FBK25]. quasi [MR25]. quasi-linear [MR25]. QUAV [ZLCL25].

radial [TNL25]. Radio [BVZ25]. random [MCHB25]. Reachable [SZY25]. Redefined [PP25]. redundant [YT25]. regularity [Tan25]. Reinforcement [LHZL25]. Relaxed [KL25]. reliability [BQS25]. Rényi [MMNP25]. replanning [KNKR25]. reputation [DWL^{+}25]. reputation-based [DWL^{+}25]. respiratory [FBK25]. responses [AS21, AS25]. results [GKM25]. retraction [CFS25]. reward [DWL^{+}25]. rewards [GLHS25]. Riemannian [CFS25]. right [CF25]. right-hand [CF25]. risk [RS25]. robust [RPS^{+}25]. rock [LDT^{+}25]. rod [MBF25]. Roman [GVW25]. rotation [ZHL25]. row [Pal25].

sampled [WSX^{+}25]. sampled-data [WSX^{+}25]. sampling [QX25]. sampling-based [QX25]. saturation [QX25]. scalar [TNL25]. scale [MAMFBM^{+}25]. scheme [MR25]. Schrödinger [LTH^{+}25]. scissors [LDT^{+}25]. search [Pal25]. Second [Tan25, FLLR25, LT25, SZY25]. second-derivative-free [LT25]. second-kind [FLLR25]. Second-order [Tan25, SZY25]. Secure [BAT25]. sensor [BAT25]. separation [WSX^{+}25]. set [SZY25]. shape [RPS^{+}25]. short [ZN25]. side [CF25]. Sierpiński [GCV25]. Signed [GVW25, LYM25, SHY25]. signless [ZW25]. simulations [MBF25]. single [MCHB25]. single- [MCHB25]. singularity [QX25]. singularity-free [QX25]. sliding [KNKR25]. snarks [ZHL25]. sojourn [MCH25]. solution [BDL25]. solutions [MBF25]. solver [CWWT25]. solvers [KBM25]. Solving [OGPFVV25, PP25]. Some [KK25b]. source [dOKA25]. spectra [ZW25]. spectral [GKM25, LZ25, TKM25]. splines [PP25]. split [LCL25]. split-star [LCL25]. Stability [TNL25, CCLL25, ZN25]. Stable [LJ25]. star [LCL25, QM25]. state [CCLL25, GPL25, LWS^{+}25]. state-constrained [GPL25]. state-dependent [CCLL25]. stochastic [DAG25, MCH25]. Stokes [AA25]. strategy [HWC25, LT25, ZJX25]. structure [KBM25, LTH^{+}25]. structure-preserving [LTH^{+}25]. subject [QX25]. sum [GPL25]. summations [KL25]. Superconvergence [GKM25]. sure [CMG25]. swaps [Ruj25]. switched [CCLL25, LWS^{+}25]. switching [MCH25]. synchronization [HWC25, SHY25, WSX^{+}25]. system [BQS25, FBK25, ZZ25a]. systems [BAT25, CGH^{+}25, CCLL25, HWC25, LHZL25, LWS^{+}25, MCH25, QX25, ZWQC25, ZLA25, ZZ25b].

taxis [Che25]. Taylor [BSG^{+}25]. their [BSG^{+}25]. three [LSQ25, LJ25]. three-dimensional [LJ25]. threshold [KK25a]. thresholds [FBK25]. Time [BQS25, ADPS25, CGH^{+}25, HWC25, IJM25, LHZL25, QX25, Tan25, ZWQC25, ZN25, ZLCL25]. Time-dependent [BQS25]. time-fractional [Tan25]. time-varying [LHZL25]. tolerant [CGH^{+}25, KNKR25]. Topological [MAMFBM^{+}25]. total [GVW25]. tracking [BAT25, ZWQC25, ZLA25]. trajectory [KNKR25, ZLA25]. transformation [SHY25]. transforms [dOKA25]. transport [CFS25]. treatment [BDL25]. trees [BVZ25, LSWZ25, YT25]. triggered [HWC25, LHZL25, QX25, ZWQC25, ZZ25a, ZLCL25]. tumor [BDL25]. tuning [LWS^{+}25]. Two [IAK25, LCL25, MR25, BVZ25, LJ25, Tan25, ZW25]. two-branch [BVZ25]. two-dimensional [LJ25]. Two-disjoint-cycle-cover [LCL25]. two-grid [Tan25]. Two-Level [IAK25, MR25]. type [BSG^{+}25, Liu25].

UAV [KNKR25]. unbounded [TKM25]. uncertain [BAT25, LWS^{+}25, XZN^{+}25]. unicyclic [MP25]. uniform [JSZ25, PP25, Tan25]. unknown [CGH^{+}25, QX25, ZZ25b, ZLCL25]. using [BSG^{+}25, LT25, MBF25, OGPFVV25].

validation [KBM25]. valuation [XZN^{+}25]. values [CF25]. variants [LT25]. varying [LHZL25]. Vector [Kov25, CFS25]. vessels [RPS^{+}25]. via [CCLL25]. volatility [Ruj25].

Volterra [Che25]. vs [DAG25].

wave [ALC^{+}25]. weak [Tan25]. wheel [ZLW25]. without [BSG^{+}25, GCV25, SHY25]. Wright [AA25]. written [XZN^{+}25].

Zero [ZJX25, GPL25]. Zero-determinant [ZJX25]. zero-sum [GPL25]. zone [LHZL25].

References

Askari:2025:SPM

[AA25] Hassan Askari and Alireza Ansari. Stokes phenomenon for the M-Wright function of order 1n. Applied Mathematics and Computation, 487 (??):??, February 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                                      /www.sciencedirect.com/
                                                      science/article/pii/S0096300324005496.

Araujo:2025:ITG

[ADPS25] Julio Araujo, Mitre C. Dourado, Fábio Protti, and Rudini Sampaio. The iteration time and the general position number in graph convexities. Applied Mathematics and Computation, 487 (??):??, February 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                                      /www.sciencedirect.com/
                                                      science/article/pii/S0096300324005459.

Alcantara:2025:NAN

[ALC^{+}25] Adriano A. Alcântara, Juan B.

Límaco, Bruno A. Carmo, Ronald R. Guardia, and Mauro A. Rincon. Numerical analysis for nonlinear wave equations with boundary conditions: Dirichlet, acoustics and impenetrability. Applied Mathematics and Computation, 484(??): ??, January 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324004703.

Anonymous:2025:Fa

[Ano25a] Anonymous. 1 February 2025. Applied Mathematics and Computation, 486 (??):??, February 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2025:Ja

[Ano25b]  Anonymous.  1 January

2025. Applied Mathematics and Computation, 484 (??):??, January 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2025:Fb

[Ano25c] Anonymous. 15 February 2025. Applied Mathematics and Computation, 487 (??):??, February 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2025:Jb

[Ano25d] Anonymous. 15 January 2025. Applied Mathematics and Computation, 485 (??):??, January 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2025:EBa

[Ano25e]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 484 (??):??, January 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324005101.

Anonymous:2025:EBb

[Ano25f]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 485 (??):??, January 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324005241.

Anonymous:2025:EBc

[Ano25g]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 486 (??):??, February 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324006015.

Anonymous:2025:EBd

[Ano25h]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 487 (??):??, February 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324006313.

Ackley:2021:LDN

[AS21] Matthew Ackley and Peter Stechlinski. Lexicographic derivatives of nonsmooth glucose-insulin kinetics under normal and artificial pancreatic responses. Applied Mathematics and Computation, 395(??):Article 125876, April 15, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300320308298.

See corrigendum [AS25].

Ackley:2025:CLD

[AS25] Matthew Ackley and Peter Stechlinski. Corrigendum to “Lexicographic derivatives of nonsmooth glucose– insulin kinetics under normal and artificial pancreatic responses” [Appl. Math. Compt. 395 (2021) 125876]. Applied Mathematics and Computation, 487(??):??, February 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324005940.

See [AS21].

Baghi:2025:SAO

[BAT25] Hamidreza Baghi, Farzaneh Abdollahi, and Heidar Ali Talebi. Secure adaptive output-feedback tracking control for a class of uncertain nonlinear systems in the presence of sensor and actuator faults under DoS and injection attacks. Applied Mathematics and Computation, 487 (??):??, February 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009630032400568X.

Bulai:2025:NSM

[BDL25] I. M. Bulai, M. C. De Bonis, and C. Laurita. Numerical solution of metastatic tumor growth models with treatment. Applied Mathematics and Computation, 484(??):??, January 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324004491.

Bilfaqih:2025:TDR

[BQS25] Yusuf Bilfaqih, Mochamad Nur Qomarudin, and Mochammad Sahal. Time-dependent reliability computation of system with multistate components. Applied Mathematics and Computation, 487 (??):??, February 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005435.

Bate:2025:JTM

[BSG^{+}25] Indra Bate, Kedarnath Senapati, Santhosh George, Muniyasamy M., and Chandhini G. Jarratt-type methods and their convergence analysis without using Taylor expansion. Applied Mathematics and Computation, 487(??):??, February 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005733.

Bantva:2025:RLT

[BVZ25] Devsi Bantva, Samir Vaidya, and Sanming Zhou. Radio labelling of two-branch trees. Applied Mathematics and Computation, 487 (??):??, February 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005587.

Cui:2025:ESS

[CCLL25] Qian Cui, Jinde Cao, Lulu Li, and Yang Liu. Exponential stability of switched systems with state-dependent delayed impulses via Bequivalent method. Applied Mathematics and Computation, 486(??):??, February 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324004922.

Cococcioni:2025:LPI

[CF25] Marco Cococcioni and Lorenzo Fiaschi. Linear programming with infinite, finite, and infinitesimal values in the right-hand side. Applied Mathematics and Computation, 486(??):??, February 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324005058.

Chen:2025:NCG

[CFS25] Kangming Chen, Ellen Hidemi Fukuda, and Hiroyuki Sato. Nonlinear conjugate gradient method for vector optimization on Riemannian manifolds with retraction and vector transport. Applied Mathematics and Computation, 486 (??):??, February 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324004624.

Cheng:2025:FTA

[CGH^{+}25] Xuanrui Cheng, Ming Gao, Wuxiang Huai, Yichun Niu, and Li Sheng. Fixedtime active fault-tolerant control for dynamical systems with intermittent faults and unknown disturbances. Applied Mathematics and Computation, 486(??):??, February 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005150.

Chen:2025:PDL

[Che25] Mengxin Chen. Pattern dynamics of a Lotka–Volterra model with taxis mechanism. Applied Mathematics and Computation, 484 (??):??, January 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324004788.

Chen:2025:MEE

[CLLC25] Xiaofeng Chen, Dongyuan Lin, Hua Li, and Zhi Cheng. Minimum error entropy high-order extend Kalman filter with fiducial points. Applied Mathematics and Computation, 487 (??):??, February 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324005745.

Canovas:2025:DLH

[CMG25] J. S. Cánovas and M. Muñoz-Guillermo. On the dynamics of a linear-hyperbolic population model with Allee effect and almost sure extinction. Applied Mathematics and Computation, 486 (??):??, February 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324004661.

Chen:2025:FBD

[CWWT25] Jianjun Chen, Yuxuan Wang, Weiyi Wang, and Zhijun Tan. A FFT-based DDSIIM solver for elliptic interface problems with discontinuous coefficients on arbitrary domains and its error analysis. Applied Mathematics and Computation, 487 (??):??, February 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324005472.

Divya:2025:MPP

[DAG25] G. Divya, S. Athithan, and Mini Ghosh. Modelling the prevalence of prostitution under the influence of poverty: a deterministic vs. stochastic approach. Applied Mathematics and Computation, 485 (??):??, January 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324004910.

deOliveira:2025:ITE

[dOKA25] André J. P. de Oliveira, Diego C. Knupp, and Luiz A. S. Abreu. Integral transforms for explicit source estimation in non-linear advection–diffusion problems. Applied Mathematics and Computation, 487 (??):??, February 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005538.

Deng:2025:RBR

[DWL^{+}25] Lili Deng, Rugen Wang, Ying Liao, Ronghua Xu, and Cheng Wang. The reputation-based reward mechanism promotes the evolution of fairness. Applied Mathematics and Computation, 486(??):??, February 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005034.

Fome:2025:ACR

[FBK25] Anna Daniel Fome, Wolfgang Bock, and Axel Klar. Analysis of a competitive respiratory disease system with quarantine: Epidemic thresholds and crossimmunity effects. Applied Mathematics and Computation, 485(??):??, January 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324004296.

Fermo:2025:GAM

[FLLR25] Luisa Fermo, Anna Lucia Laguardia, Concetta Laurita, and Maria Grazia Russo. A global approximation method for secondkind nonlinear integral equations. Applied Mathematics and Computation, 487 (??):??, February 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324005551.

Gurubachan:2025:AFF

[GCV25] Gurubachan, V. V. M. S. Chandramouli, and S. Verma. Analysis of α-fractal functions without boundary point conditions on the Sierpiński gasket. Applied Mathematics and Computation, 486(??):??, February 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005332.

Gupta:2025:SRH

[GKM25] Saloni Gupta, Arnab Kayal, and Moumita Mandal. Superconvergence results for hypersingular integral equation of first kind by Chebyshev spectral projection methods. Applied Mathematics and Computation, 487(??):??, February 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630032400554X.

Geng:2025:ECA

[GLHS25] Yini Geng, Yikang Lu, Lijun Hong, and Lei Shi. Evolution of cooperation with asymmetric rewards. Applied Mathematics and Computation, 487(??):??, February 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005368.

Guo:2025:DED

[GPL25] Siyu Guo, Yingnan Pan, and Hongyi Li. Dynamic event-driven optimal consensus control for stateconstrained multiagent zerosum differential graphical games. Applied Mathematics and Computation, 484 (??):??, January 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324004405.

Guo:2025:STR

[GVW25] Yubao Guo, Lutz Volkmann, and Yun Wang. Signed total Roman domination and domatic numbers in graphs. Applied Mathematics and Computation, 487 (??):??, February 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324005356.

Hou:2025:PFT

[HWC25] Xinzhao Hou, Huaiqin Wu, and Jinde Cao. Practical finite-time synchronization for Lur’e systems with performance constraint and actuator faults: a memorybased quantized dynamic event-triggered control strategy. Applied Mathematics and Computation, 487 (??):??, February 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324005691.

Iqbal:2025:TLM

[IAK25] Azhar Iqbal, Shahbaz Ahmad, and Junseok Kim. Two-level method for blind image deblurring problems. Applied Mathematics and Computation, 485(??):??, January 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324004697.

Iman:2025:DDP

[IJM25] Sara Iman and Mohammad-Reza Jahed-Motlagh. Datadriven predictor of controlaffine nonlinear dynamics: Finite discrete-time bilinear approximation of Koopman operator. Applied Mathematics and Computation, 487(??):??, February 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005290.

Jiang:2025:NFN

[JSZ25] Taijiang Jiang, Qiang Sun, and Chao Zhang. A note on the 1-factorization of nonuniform complete hypergraph. Applied Mathematics and Computation, 485 (??):??, January 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324004685.

Kunstek:2025:MCV

[KBM25] Petar Kunstek, Martina Bukac, and Boris Muha. Mass conservation in the validation of fluid-poroelastic structure interaction solvers. Applied Mathematics and Computation, 487(??):??, February 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324005423.

Keng:2025:CPL

[KK25a] Ying Ying Keng and Kiam Heong Kwa. Contagion probability in linear threshold model. Applied Mathematics and Computation, 487 (??):??, February 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324005514.

Kim:2025:SID

[KK25b] Taekyun Kim and Dae San Kim. Some identities on degenerate harmonic and degenerate higher-order harmonic numbers. Applied Mathematics and Computation, 486(??):??, February 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009630032400506X.

Kim:2025:RCP

[KL25] Do Wan Kim and Donghwan Lee. Relaxed conditions for parameterized linear matrix inequality in the form of nested fuzzy summations. Applied Mathematics and Computation, 487 (??):??, February 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009630032400540X.

Khodaverdian:2025:FTM

[KNKR25] Maria Khodaverdian, Majdeddin Najafi, Omid Kazemifar, and Shahabuddin Rahmanian. Fault-tolerant model predictive sliding mode control for trajectory replanning of multi-UAV formation flight. Applied Mathematics and Computation, 487(??):??, February 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005344.

Kovacevic:2025:VMM

[Kov25] Mladen Kovacević. Vector multispaces and multispace codes. Applied Mathematics and Computation, 486 (??):??, February 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324005022.

Li:2025:TDC

[LCL25] Hao Li, Liting Chen, and Mei Lu. Two-disjoint-cycle-cover pancyclicity of splitstar networks. Applied Mathematics and Computation, 487(??):??, February 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324005460.

Lu:2025:EPR

[LDT^{+}25] Yikang Lu, Hui Dai, Huaiyu Tan, Xiaofang Duan, Lei Shi, and Junpyo Park. Enhancement of persistence in the rock–paper–scissors dynamics through higherorder interactions. Applied Mathematics and Computation, 487(??):??, February 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324005447.

Li:2025:RLB

[LHZL25] Xin Li, Dakuo He, Qiang Zhang, and Hailong Liu. Reinforcement learning-based adaptive event-triggered control of multi-agent systems with time-varying deadzone. Applied Mathematics and Computation, 486 (??):??, February 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005204.

Liu:2025:ECN

[Liu25] Qiong Liu. The equivalent conditions for norm of a Hilbert-type integral operator with a combination kernel and its applications. Applied Mathematics and Computation, 487 (??):??, February 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009630032400537X.

Li:2025:SGF

[LJ25] Jiajun Li and Ying Jiang.

Stable generalized finite element for two-dimensional and three-dimensional nonhomogeneous interface problems. Applied Mathematics and Computation, 486 (??):??, February 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005174.

Liu:2025:PEP

[LL25] Yandi Liu and Yonghui Li. A payoff equality perspective for evolutionary games: Mental accounting and cooperation promotion. Applied Mathematics and Computation, 486 (??):??, February 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324005009.

Lotstedt:2025:NIM

[LM25] Per Lötstedt and Sonja Mathias. Numerical integration of mechanical forces in center-based models for biological cell populations. Applied Mathematics and Computation, 486(??):??, February 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324005307.

Li:2025:GBN

[LSQ25] Yinkui Li, Guiyu Shi, and Xiaoxiao Qin. Graphs with burning number three. Applied Mathematics and Computation, 487(??):??, February 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324005617.

Liu:2025:NCE

[LSWZ25] Fenjin Liu, Ke Su, Wei Wang, and Hao Zhang. Non-cospectral equienergetic trees of diameter at most four. Applied Mathematics and Computation, 487(??):??, February 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005654.

Laadhari:2025:EFE

[LT25] Aymen Laadhari and Helmi Temimi. Efficient finite element strategy using enhanced high-order and second-derivative-free variants of Newton’s method. Applied Mathematics and Computation, 486(??):??, February 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005198.

Luo:2025:LSP

[LTH^{+}25] Fangwen Luo, Qiong Tang, Yiting Huang, Yanhui Ding, and Sijia Tang. Local structure-preserving algorithms for the nonlinear Schrödinger equation with power law nonlinearity. Applied Mathematics and Computation, 484(??): ??, January 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324004478.

Liu:2025:DOB

[LWS^{+}25] Ming-Rui Liu, Li-Bing Wu, Hong Sang, Liang-Dong Guo, and Sheng-Juan Huang. Disturbance observer-based fuzzy adaptive switched tuning control of uncertain nonlinear systems with full state constraints. Applied Mathematics and Computation, 487 (??):??, February 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009630032400571X.

Li:2025:MLL

[LYM25] Dan Li, Minghui Yan, and Jixiang Meng. Minimizing the least Laplacian eigenvalue of signed complete graphs. Applied Mathematics and Computation, 484 (??):??, January 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324004636.

Li:2025:SEG

[LZ25] Shuchao Li and Zihan Zhou. On spectral extrema of graphs with given order and generalized 4-independence number. Applied Mathematics and Computation, 484 (??):??, January 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009630032400479X.

Molina-Abril:2025:TSF

[MAMFBM^{+}25] H. Molina-Abril, M. J.

Morón-Fernández, M. Benito-Marimón, F. Díaz del Río, and P. Real. Topological scale framework for hypergraphs. Applied Mathematics and Computation, 485 (??):??, January 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324004508.

Meier:2025:MFS

[MBF25] Andreas Meier, Eberhard Bänsch, and Florian Frank. Multiscale flow simulations of dilute polymeric solutions with bead-rod chains using Brownian configuration fields. Applied Mathematics and Computation, 487 (??):??, February 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005526.

Mei:2025:ACA

[MCH25] Xu Mei, Jun Cheng, and Wentao Huang. Attackcompensated asynchronous output feedback control for stochastic switching systems with sojourn probability. Applied Mathematics and Computation, 485 (??):??, January 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324004855.

Ma:2025:IAO

[MCHB25] Liyang Ma, Daniel Conus, Wei-Min Huang, and Paolo Bocchini. Improved algorithm for the optimal quantization of single- and multivariate random functions. Applied Mathematics and Computation, 486 (??):??, February 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324004892.

Munoz:2025:ELL

[MMNP25] Francisco J. Muñoz, Luca Meacci, Juan Carlos Nuño, and Mario Primicerio. Exploring the limits of the law of mass action in the mean field description of epidemics on Erdős–Rényi networks. Applied Mathematics and Computation, 485 (??):??, January 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324004806.

Moon:2025:LED

[MP25] Sunyo Moon and Seungkook Park. Laplacian eigenvalue distribution for unicyclic graphs. Applied Mathematics and Computation, 485 (??):??, January 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324004831.

Mittal:2025:TLI

[MR25] Kajal Mittal and Rajendra K. Ray. Two-level implicit high-order compact scheme in exponential form for 3D quasilinear parabolic equations. Applied Mathematics and Computation, 485(??):??, January 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324004454.

Orizaga:2025:SAC

[OGPFVV25] Saulo Orizaga, Gilberto González-Parra, Logan Forman, and Jesus Villegas-Villanueva. Solving Allen– Cahn equations with periodic and nonperiodic boundary conditions using mimetic finite-difference operators. Applied Mathematics and Computation, 484(??):??, January 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324004545.

Palubeckis:2025:FLS

[Pal25] Gintaras Palubeckis. A fast local search based memetic algorithm for the parallel row ordering problem. Applied Mathematics and Computation, 486(??):??, February 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324005010.

Palav:2025:RFO

[PP25] Mansi S. Palav and Vikas H. Pradhan. Redefined fourth order uniform hyperbolic polynomial B-splines based collocation method for solving advection-diffusion equation. Applied Mathematics and Computation, 484 (??):??, January 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324004533.

Qiao:2025:DPC

[QM25] Hongwei Qiao and Jixiang Meng. Disjoint path covers of star graphs. Applied Mathematics and Computation, 487(??):??, February 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005599.

Qi:2025:ETS

[QX25] Xiaojing Qi and Shengyuan Xu. Event-triggered samplingbased singularity-free fixedtime control for nonlinear systems subject to input saturation and unknown control directions. Applied Mathematics and Computation, 486(??):??, February 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005319.

Romero:2025:RSM

[RPS^{+}25] Pau Romero, Abel Pedrós, Rafael Sebastian, Miguel Lozano, and Ignacio García-Fernández. A robust shape model for blood vessels analysis. Applied Mathematics and Computation, 487 (??):??, February 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005393.

Ramirez:2025:OII

[RS25] Hugo E. Ramírez and Rafael Serrano. Optimal investment with insurable background risk and nonlinear portfolio allocation frictions. Applied Mathematics and Computation, 485 (??):??, January 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324004843.

Rujivan:2025:APV

[Ruj25] Sanae Rujivan. Analytically pricing volatility options and capped/floored volatility swaps with nonlinear payoffs in discrete observation case under the Merton jumpdiffusion model driven by a nonhomogeneous Poisson process. Applied Mathematics and Computation, 486 (??):??, February 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324004909.

Sun:2025:PPB

[SHY25] Yu Sun, Cheng Hu, and Juan Yu. Pinning passivity and bipartite synchronization of fractional signed networks without gauge transformation. Applied Mathematics and Computation, 486(??):??, February 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324005289.

Suvetha:2025:NFO

[SNP25] R. Suvetha, J. J. Nieto, and P. Prakash. Nonfragile output-feedback control for delayed memristive bidirectional associative memory neural networks against actuator failure. Applied Mathematics and Computation, 485 (??):??, January 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009630032400482X.

Shen:2025:RSE

[SZY25] Yi Shen, Jiemei Zhao, and Liqi Yu. Reachable set estimation of delayed secondorder memristive neural networks. Applied Mathematics and Computation, 484 (??):??, January 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324004557.

Tan:2025:SON

[Tan25] Zhijun Tan. Second-order non-uniform and fast twogrid finite element methods for non-linear timefractional mobile/immobile equations with weak regularity. Applied Mathematics and Computation, 486 (??):??, February 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324005046.

Tissaoui:2025:ESE

[TKM25] Yassine Tissaoui, James F. Kelly, and Simone Marras. Efficient spectral element method for the Euler equations on unbounded domains. Applied Mathematics and Computation, 487 (??):??, February 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324005411.

Tominec:2025:SER

[TNL25] Igor Tominec, Murtazo Nazarov, and Elisabeth Larsson. Stability estimates for radial basis function methods applied to linear scalar conservation laws. Applied Mathematics and Computation, 485 (??):??, January 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324004818.

Wang:2025:MSB

[WSX^{+}25] H.-Z. Wang, X.-C. Shangguan, D. Xiong, Y.-H. An, and L. Jin. A matrix-separation-based integral inequality for aperiodic sampled-data synchronization of delayed neural networks considering communication delay. Applied Mathematics and Computation, 486(??):??, February 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324004934.

Xin:2025:EPV

[XZN^{+}25] Yue Xin, Yi Zhang, Idin Noorani, Farshid Mehrdoust, and Jinwu Gao. Estimation of parameters and valuation of options written on multiple assets described by uncertain fractional differential equations. Applied Mathematics and Computation, 487(??):??, February 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005708.

Yang:2025:ERT

[YT25] Qing Yang and Yingzhi Tian. Exploring redundant trees in bipartite graphs. Applied Mathematics and Computation, 486(??):??, February 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324004673.

You:2025:GNH

[YYW^{+}25] Tao You, Linjiang Yang, Jian Wang, Peng Zhang, Jinchao Chen, and Ying Zhang. The guidance of neutral human populations maintains cooperation in the prisoner’s dilemma game. Applied Mathematics and Computation, 486(??):??, February 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005320.

Zhou:2025:PMC

[ZHL25] Wenjuan Zhou, Rong-Xia Hao, and Yilun Luo. Perfect matching cover indices of generalized rotation snarks. Applied Mathematics and Computation, 487 (??):??, February 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005629.

Zhang:2025:ZDS

[ZJX25] Zhipeng Zhang, Xiaotong Jiang, and Chengyi Xia. Zero-determinant strategy of finite games with implementation errors and its application into group decisionmaking. Applied Mathematics and Computation, 486 (??):??, February 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005162.

Zhang:2025:BFB

[ZLA25] Lili Zhang, Chenglong Liu, and Liwei An. Barrierfunction based adaptive trajectory tracking control for high-order nonlinear systems with collision avoidance. Applied Mathematics and Computation, 484 (??):??, January 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009630032400465X.

Zhu:2025:DET

[ZLCL25] Guanchao Zhu, Min Luo, Guozeng Cui, and Ze Li. Dynamic event-triggered predefined-time adaptive attitude control for a QUAV with unknown deception attacks. Applied Mathematics and Computation, 485 (??):??, January 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324004648.

Zhang:2025:CCW

[ZLW25] Guozhen Zhang, Xin Liu, and Dajin Wang. Component connectivity of wheel networks. Applied Mathematics and Computation, 487(??):??, February 15, 2025. CODEN AMHCBQ.

ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300324005575.

Zhang:2025:MES

[ZN25] Jinsen Zhang and Xiaobing Nie. Multiple exponential stability for short memory fractional impulsive Cohen–Grossberg neural networks with time delays. Applied Mathematics and Computation, 486 (??):??, February 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324005277.

Zhang:2025:DSL

[ZW25] Yonghong Zhang and Ligong Wang. Distance (signless) Laplacian spectra and energies of two classes of cyclic polyomino chains. Applied Mathematics and Computation, 487(??):??, February 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324005605.

Zhai:2025:EBE

[ZWQC25] Junchang Zhai, Huanqing Wang, Yuping Qin, and Hongxia Cui. Estimationbased event-triggered fixedtime fuzzy tracking control for high-order nonlinear systems. Applied Mathematics and Computation, 486 (??):??, February 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005186.

Zheng:2025:ETD

[ZZ25a] Leyi Zheng and Yimin Zhou.

Event-triggered distributed consensus control of heterogeneous multi-agent system under communication and actuator faults. Applied Mathematics and Computation, 487(??):??, February 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005381.

Zhou:2025:ADU

[ZZ25b] Dan-Dan Zhou and Ran Zhao. Adaptive distributed unknown input observer for linear systems. Applied Mathematics and Computation, 486(??):??, February 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324004880.